### v1

# Last Update: Steptember 24, 2025

Last Update: Steptember 24, 2025
On Structured State-Space Duality
Jerry Yao-Chieh Hu†‡∗111jhu@u.northwestern.edu; jhu@ensemblecore.ai  Xiwen Zhang§∗222xiz284@ucsd.edu  Weimin Wu333wwm@u.northwestern.edu  Han Liu†♯444hanliu@northwestern.edu
**footnotetext: These authors contributed equally to this work. Part of the work done during JH’s internship at Ensemble AI.
[…]
Department of Computer Science, Northwestern University, Evanston, IL 60208, USA
[…]
Structured State-Space Duality (SSD) [Dao & Gu, ICML 2024] is an equivalence between a simple Structured State-Space Model (SSM) and a masked attention mechanism. In particular, a state-space model with a scalar-times-identity state matrix is equivalent to a masked self-attention with a $1$-semiseparable causal mask. Consequently, the same sequence transformation (model) has two algorithmic realizations: as a linear-time $O(T)$ recurrence or as a quadratic-time $O(T^{2})$ attention. In this note, we formalize and generalize this duality: (i) we extend SSD from the scalar‑identity case to general diagonal SSMs (diagonal state matrices); (ii) we show that these diagonal SSMs match the scalar case’s training complexity lower bounds while supporting richer dynamics; (iii) we establish a necessary and sufficient condition under which an SSM is equivalent to $1$-semiseparable masked attention; and (iv) we show that such duality fails to extend to standard softmax attention due to rank explosion. Together, these results tighten bridge between recurrent SSMs and Transformers, and widen the design space for expressive yet efficient sequence models.
Contents
1 Introduction
Structured State-Space Duality (SSD) refers to a one-to-one equivalence between a certain linear Structured State-Space Model (SSM) and a masked self-attention mechanism (Dao and Gu, 2024). In plain terms, it means the same sequence transformation has two algorithmic realizations: either as a recurrent state-space system or as a self-attention (matrix) operation. Dao and Gu (2024) introduce the first example of such duality: a state-space model whose state matrix is a scalar multiple of the identity is equivalent to a causal self-attention with a rank-1 mask matrix.
In this case, we write the sequence (time) index of a matrix as a superscript ($A^{t}$). Then the SSM update $h_{t+1}=A^{t}h_{t}+b_{t}x_{t}$ (with $h_{1}=0$ and output $y_{t}=c_{t}^{\top}h_{t}$ for t=1,,Tt=1,...,T) yields the closed-form solution $y_{t}=\sum_{s=1}^{t}c_{t}^{\top}A^{t}\cdots A^{s+1}b_{s}x_{s}$. Equivalently, the self-attention viewpoint treats $y$ as an explicit attention matrix $M\in\mathbb{R}^{T\times T}$ acting on $x$, with entries $M_{t,s}=c_{t}^{\top}A^{t}\cdots A^{s+1}b_{t}$ for $s\leq t$ (and $M_{t,s}=0$ for $s>t$ due to the causal mask). This attention matrix $M$ is a rank-$N$ matrix with a $1$-semiseparable (Definition˜3.2) causal mask, where $N$ is the dimension of $A^{t}$, also known as the state dimension. Thus, the SSM and the masked attention realize the same function $x\mapsto y$: one via a linear-time $O(T)$ recurrence and the other via a quadratic-time $O(T^{2})$ matrix multiplication.
Remarkably, this duality bridges two disparate paradigms for sequence modeling — recurrent state-space models (Dao and Gu, 2024; Gu and Dao, 2023) and Transformer attention (Vaswani et al., 2017). State-space models update a latent state recurrently, and hence yields linear complexity in sequence length. Attention mechanisms compute pairwise token interactions, and hence yields quadratic complexity. The structured state-space duality unifies these two paradigms by revealing that they implement identical functions.
Nevertheless, while Dao and Gu (2024) conjecture that analogues duality should hold for diagonal state-space models, there exist no formal treatment to the best of our knowledge. We give a concrete diagonal state-space duality, and provide the regarding computation algorithm.
Contributions.
In this work, we build on SSD and extend its scope in four key directions:
  • General Diagonal SSMs (Sections˜4.1 and 4.2). We extend SSD beyond the simple $(\text{scalar})\times I_{N}$ state matrix to general diagonal state matrices. This enlarges the class of SSMs under the duality (from a single exponential decay to $N$ separate diagonal dynamics), thereby supporting richer sequence dynamics.
  • Efficiency at Scale (Section˜4.3). We prove that these more general diagonal SSMs match the training complexity lower bound of the scalar case while offering greater expressiveness. In other words, it is possible to train and execute the richer diagonal SSMs with the same optimal $O(TN)$ time complexity as the scalar SSM, so we get additional modeling power at no extra asymptotic cost.
  • Higher-Rank Equivalence (Section˜A.1). We prove the conjecture of (Dao and Gu, 2024) that the equivalence between semiseparable matrices and matrices with sequential state-space representation holds not only for the rank-1 case, but also for general rank $N$. In other words, an SSM of state dimension $N$ corresponds to an $N$-semiseparable attention matrix.
  • Masked Attention Duality of General SSM (Section˜4.4). While each $1$-semiseparable masked attention has an SSM dual, we provide a necessary and sufficient condition for an $N$-semiseparable (Definition˜3.1) matrix (corresponding to an SSM of state dimension $N$) to have a $1$-semiseparable masked attention dual.
Together, these results strengthen the bridge between recurrent state-space models and Transformer-style attention mechanism. They also widen the design space for expressive yet efficient sequence modeling. Through this generalized duality framework, we enable principled exploration of new architectures that enjoy the best of both worlds (recurrent and attention).
Organization.
We provide related work in Section˜2, and background in Section˜3. We show our main theory in Section˜4, and the limitations of structured state-space duality in Section˜5.
Notations.
We denote the index set $\{1,\ldots,I\}$ by $[I]$. We denote vectors with lower case and matrices with upper case. We write the sequence (time) index of a matrix as a superscript ($A^{t}$). We write the sequence (time) index of a vector or scalar as a subscript ($a_{t}$). For a matrix $A$, $A_{i,j}$ denotes its entry on the $i$-th row and the $j$-th column. $A_{i,:}$ denotes its $i$-th row, $A_{:,j}$ denotes its $j$-th column. $A_{i_{1}:i_{2},j_{1}:j_{2}}\in\mathbb{R}^{(i_{2}-i_{1})\times(j_{2}-j_{1})}$ denotes its submatrix containing all its entries $A_{i,j}$ where $i_{1}\leq i\leq i_{2}-1$ and $j_{1}\leq j\leq j_{2}-1$. We use $\odot$ for Hadamard multiplication. We write the input sequence of length $T$ as $x_{1}^{\top},x_{2}^{\top},\dots,x_{T}^{\top}]\in\mathbb{R}^{d\times T}$ and the output at time $t\in[T]$ as $y_{t}\in\mathbb{R}^{1\times d}$.
2 Related Work
In this section, we review related work on structured state-space models, efficient Transformers and linear attention mechanisms, and the connections between state-space models and attention.
Structured State-Space Models (SSMs).
SSMs aim to model long-range dependencies with linear-time computation. S3 first introduced the idea of parameterizing state-space models with structured matrices to enable efficient sequence modeling (Gu et al., 2021a). S4 introduces a state-space layer with stable diagonalization and fast convolution, which enabled long-context training and inference (Gu et al., 2021b). S5 simplifies the design with a single multi-input multi-output SSM and preserved $O(T)$ scaling (Smith et al., 2023). Mamba adds input-dependent gating on the SSM projections and achieved strong accuracy with linear-time sequence modeling (Gu and Dao, 2023). These works establish SSMs as competitive sequence models and motivate analyses that compare their expressive power to attention.
Efficient Transformers and Linear Attention.
Many works reduce the quadratic cost of self-attention by imposing structure or approximation (Tay et al., 2022). Linformer projects keys and values to low rank and reduced compute and memory while preserving accuracy (Wang et al., 2020). Linear Transformers replace softmax by kernel feature maps and execute attention as a recurrence, which yielded $O(T)$ autoregressive inference (Katharopoulos et al., 2020). Performer uses random features to approximate softmax attention with variance control and linear complexity (Choromanski et al., 2021). Nyströmformer applies Nyström approximation to attention and obtains sub-quadratic cost (Xiong et al., 2021). Sparse patterns such as Longformer, BigBird, and Reformer improve scaling by local windows, global tokens, or LSH-based routing (Beltagy et al., 2020; Zaheer et al., 2020; Kitaev et al., 2020). Retentive networks propose a recurrent retention operator that matches Transformer quality with linear-time execution (Sun et al., 2023). These methods show that attention admits efficient surrogates when the attention matrix has a low-rank, kernel, or sparse structure.
Connections between SSMs and Attention.
Linear attention admits a recurrent implementation and thus links attention with RNN-style computation (Katharopoulos et al., 2020). Dao and Gu (2024) introduce Structured State Duality (SSD), prove the duality between scalar-identity SSMs and masked attention with $1$-semiseparable kernels, and conjectrue such duality hold in diagonal SSM. Hydra generalizes matrix mixers beyond causal SSMs with quasi-separable structure and bidirectional information flow (Hwang et al., 2024).
Our work differs in scope and goal: we give an algebraic duality between $N$-dimensional diagonal SSM and $1$-semiseparable masked attention, and we prove that diagonal SSMs match the scalar case in training FLOPs and memory while enabling $N$ independent state modes. These results place diagonal SSD on a firm foundation and clarifies how SSM capacity aligns with semiseparable rank.
3 Background
In this section, we present the ideas we build on. Then we review the prior results of state-space duality (SSD) from (Dao and Gu, 2024) for the structured state-space duality in Section˜3.2.
3.1 Semiseparable Matrices
To prepare our theory, we begin by defining the class of semiseparable (SS) matrices.
NN-Semiseparable (NN-$S$) Matrix.
We give a definition of this matrix whose lower-triangular submatrices have union-bounded rank as follows.
Definition 3.1 ($N$-Semiseparable (NN-$S$) Matrix).
A lower triangular matrix $M$ is $N$-semiseparable (N-SS) if every submatrix of $M$ consisting of entries on or below the main diagonal has rank at most $N$. The smallest such $N$ is called the semiseparable rank (or order) of $M$.
In essence, an $N$-semiseparable matrix allows up to $N$ independent directions in each lower-triangular block.
11-Semiseparable (11-$S$ Matrix).
Now we define a special case of $N$-$S$ matrix with very low rank: $1$-semiseparable matrix.
Definition 3.2 ($1$-Semiseparable (11-$S$) Matrix).
Suppose $M$ is a lower triangular matrix. $M$ is 1-semiseparable ($1$-$S$) if and only if every submatrix of $M$ consisting of entries on or below the main diagonal has rank at most $1$.
Intuitively, a 1-SS matrix has extremely low complexity: each new row introduces at most one new independent direction in the space of lower-triangular entries.
Remark 3.1 (Prior Work).
We remark that, 1-$S$ is the common setup when considering $S$ matrix analysis (Fasino and Gemignani, 2002; Vandebril et al., 2005, 2008). In contrast, we consider $S$ with higher rank.
We next define a $1$-$S$ masked attention operator. In essence, it is an attention layer with $1$-$S$ matrix as its causal mask.
Definition 3.3 ($1$-$S$ Masked Attention).
Let $Q,K\in\mathbb{R}^{T\times N}$ be query and key matrices. A 1-SS masked attention is a self-attention operation whose attention weight matrix is masked by a 1-SS matrix $L\in\mathbb{R}^{T\times T}$. In particular, the attention scores take the form $L\odot(QK^{\top})$, i.e. the element-wise product of $QK^{\top}$ with the mask $L$ (with $L$ enforcing a causal lower-triangular structure).
Furthermore, we introduce the notion of a Sequentially SemiSeparable (SSS) representation. Importantly, SSS connects these structured matrices to state-space models:
Definition 3.4 ($N$-Sequentially Semiseparable ($N$-SSS) Representation).
A lower triangular matrix $M\in\mathbb{R}^{T\times T}$ has an $N$-sequentially semiseparable (N-SSS) representation if there exist vectors $b_{1},\dots,b_{T}\in\mathbb{R}^{N}$, $c_{1},\dots,c_{T}\in\mathbb{R}^{N}$, and matrices $A^{1},\dots,A^{T}\in\mathbb{R}^{N\times N}$ such that
$M_{j,i}=c_{j}^{\top}A^{j}\cdots A^{i+1}b_{i},$ (3.1)
for all $1\leq i\leq j\leq T$.
Definition 3.5 ($N$-SSS Representable Matrix).
A matrix $M$ is $N$-SSS representable if it admits an $N$-SSS representation (i.e., (3.1)). Equivalently, $M$ can be written in the form of (3.1).
The above definitions formalize how a structured state-space model of dimension $N$ gives rise to a matrix $M$ with semiseparable rank $N$. In particular, any $M$ that has an $N$-SSS representation is necessarily $N$-semiseparable (since each new state dimension contributes at most one new rank to the growing matrix). We next review the known correspondence between such structured matrices and attention mechanisms in the simplest (rank-1) case.
3.2 Structured State-Space Duality (SSD)
We now describe the structured state-space duality as introduced by Dao and Gu (2024) for the scalar-identity state-space case. We begin by formulating the state-space model and its induced sequence kernel, then show how it corresponds to a masked attention operator.
Time-Varying SSM and Induced Kernel.
Consider a time-varying linear state-space model (SSM) with state dimension $N$, defined by the recurrence
$\underbrace{h_{t}}_{N\times d}:=\underbrace{A^{t}}_{N\times N}\underbrace{h_{t-1}}_{N\times d}+\underbrace{b_{t}}_{N\times 1}\underbrace{x_{t}}_{1\times d},\quad\underbrace{y_{t}}_{1\times d}=\underbrace{c_{t}^{\top}}_{1\times N}\underbrace{h_{t}}_{N\times d},\quad\text{for}\quad t\in[T],$ (3.2)
where we set $h_{0}=0$ for consistency. Here, $h_{t}\in\mathbb{R}^{N\times d}$ is the hidden state, $A^{t}\in\mathbb{R}^{N\times N}$ is the state transition matrix, $b_{t}\in\mathbb{R}^{N\times 1}$ and $c_{t}\in\mathbb{R}^{N\times 1}$ are input and output weight matrices. Importantly, this recurrence defines a causal linear operator on the input sequence. Unrolling the recurrence yields an explicit input-output relation
$y_{t}$ $=\sum_{s=1}^{t}M_{t,s}x_{s},\quad\text{where}\quad M_{t,s}:=\begin{cases}c_{t}^{\top}A^{t}\cdots A^{s+1}b_{s},&\quad\text{for}\quad t\geq s;\\ 0,&\quad\text{for}\quad t<s.\end{cases}$ (3.3)
for $1\leq s\leq t\leq T$. We refer to $M_{t,s}$ as the SSM kernel at $(t,s)$. Let $M\in\mathbb{R}^{T\times T}$ denote the lower-triangular matrix of kernel coefficients, i.e. $M_{t,s}$ for $t\geq s$ (and $M_{t,s}=0$ for $t<s$). By construction, $M$ encodes the entire transformation from inputs $x_{1},\dots,x_{T}$ to outputs $y_{1},\dots,y_{T}$. Moreover, $M$ is structured: since the latent state is $N$-dimensional, $M$ has semiseparable rank at most $N$ (each row of $M$ lies in an $N$-dimensional subspace). In particular, $M$ is an $N$-$S$ matrix in the sense of Definition˜3.1, and for this special case $N=1$, $M$ is $1$-$S$.
Scalar-Times-Identity State Matrix ($A^{t}=a_{t}I_{N}$). A simple case of the above is when each state matrix is a scalar multiple of the identity. We call such an SSM a scalar-identity SSM, meaning $A_{t}=a_{t}I_{N}$ for some scalar $a_{t}\in\mathbb{R}$. In this case, the recurrence (3.2) simplifies to
$y_{t}=\sum_{s=1}^{t}a_{t}\cdots a_{s+1}c_{t}^{\top}b_{s}x_{s},$ (3.4)
which is a convolution-style sum over past inputs. For example, if $a_{t}=a$ is constant, then (3.4) reduces to the standard discrete-time convolution $y_{t}=\sum_{s=1}^{t}a^{t-s}c_{t}^{\top}b_{s}x_{s}$.
Scalar-Identity SSM.
We call an SSM layer a scalar-identity SSM if each of the state matrices $A^{t}$ is a scale multiple of the identity matrix.
Rank-11 Special Case ($N=1$).
In the extreme case of state dimension $N=1$, the state $h_{t}$ is one-dimensional. Then $b_{t}$ and $c_{t}$ are scalars for all $t$. We collect the input sequence into a matrix $X=[x_{1};x_{2};\dots;x_{T}]\in\mathbb{R}^{T\times d}$ (with $x_{t}^{\top}$ as the $t$-th row) and similarly $Y=[y_{1};\dots;y_{T}]\in\mathbb{R}^{T\times d}$ for the outputs. Let $p=(b_{1},\dots,b_{T})^{\top}\in\mathbb{R}^{T}$ and $q=(c_{1},\dots,c_{T})^{\top}\in\mathbb{R}^{T}$ denote the vectors of input and output weights over time. The scalar-identity formula (3.4) then reduces to
$\underbrace{Y}_{T\times d}=\underbrace{\text{diag}(p)}_{T\times T}\underbrace{M}_{T\times T}\underbrace{\text{diag}(q)}_{T\times T}\underbrace{X}_{T\times d},$
where
$M_{t,s}:=\begin{cases}a_{t}\cdots a_{s+1},&\quad\text{for}\quad t\geq s;\\ 0,&\quad\text{for}\quad t<s.\end{cases}$
Here $M$ is a 1-semiseparable mask matrix (Definition 3.2). In other words, the sequence mapping implemented by this $N=1$ SSM can be viewed as a masked attention operation: $M$ serves as a causal mask on the outer-product matrix $CB^{\top}=qp^{\top}$. In the notation of Definition˜3.3, this is a $1$-$S$ masked attention with $Q=C$ and $K=B$.
Now we are ready to present the structured state-space duality by Dao and Gu (2024):
Proposition 3.1 (Dao and Gu (2024) Scalar-Identity State-Space Duality).
Consider the SSM defined by (3.2) where each $A^{t}=a_{t}I_{N}$ (i.e. a scalar-identity SSM). Let $B=[b_{1};b_{2};\dots;b_{T}]^{\top}\in\mathbb{R}^{T\times N}$ and $C=[c_{1};c_{2};\dots;c_{T}]^{\top}\in\mathbb{R}^{T\times N}$ be the matrices whose $t$-th rows are $b_{t}^{\top}$ and $c_{t}^{\top}$, respectively. Define $M\in\mathbb{R}^{T\times T}$ by $M_{t,s}=a_{t}a_{t-1}\cdots a_{s+1}$ for $t\geq s$ and $M_{t,s}=0$ for $t<s$. Then for any input sequence $X=[x_{1};\dots;x_{T}]\in\mathbb{R}^{T\times d}$ with output $Y=[y_{1};\dots;y_{T}]\in\mathbb{R}^{T\times d}$, the recurrence (3.4) is equivalent to a $1$-$S$ masked attention representation:
$Y=(M\odot(CB^{\top}))X,$
where
$M_{t,s}:=\begin{cases}a_{t}\cdots a_{s+1},&\quad\text{for}\quad t\geq s;\\ 0,&\quad\text{for}\quad t<s.\end{cases}$
Here $\odot$ denotes elementwise (Hadamard) product. In particular, the same sequence transformation is realizable either by the linear-time recurrence (3.2) or by the quadratic-time matrix operation on the right-hand side.
Proposition˜3.1 (from Dao and Gu (2024)) establishes a one-to-one correspondence between a simple structured SSM and a masked self-attention operator with a 1-SS (rank-1) mask.
4 Main Theory
This section presents our main results: (i) we extend the state-space duality to general diagonal SSMs in Section˜4.1; (ii) we specify the case of state-space duality for diagonal SSMs with full-rank state matrices in Section˜4.2; (iii) we give a concrete algorithm for computing diagonal SSD layer and show that the computational complexity of diagonal SSD is the same as scalar-identity SSD in Section˜4.3; and (iv) we find the exact class of general SSMs that have 1-SS masked attention dual in Section˜4.4.
4.1 Structured State-Space Duality for General Diagonal SSMs
Dao and Gu (2024) study state-space duality of SSM with scalar-identity state matrices. Here we complement (Dao and Gu, 2024) by extending state-space duality to SSM with general diagonal state matrices. In the case of general diagonal SSM, each state matrix $A^{t}$ is a diagonal matrix. The diagonal state-space model has an attention-like dual as the sum of $N$ attention matrices.
Attention-Like Dual of Diagonal SSMs.
Suppose $M\in\mathbb{R}^{T\times T}$ is a lower triangular matrix as in (3.1) regarding the state-space model. We show that $M$ has an attention-like dual as the sum of $N$ attention-like matrices $M^{n}=L^{n}\odot(Q^{n}\cdot K^{n\top})$, where for all $n\in[N]$ we have $Q^{n},K^{n}\in\mathbb{R}^{T\times 1}$.
Figure 1: $M_{j,i}=c_{j}^{\top}A^{j}\cdots A^{i+1}b_{i}$
Figure 2: Construction of $b^{n}$ and $c^{n}$.
Specifically, suppose
$M_{j,i}=c_{j}^{\top}A^{j}\cdots A^{i+1}b_{i},\quad\text{for all}\quad 1\leq i\leq j\leq T,$ (4.1)
where $b_{1},\cdots,b_{T},c_{1},\cdots,c_{T}\in\mathbb{R}^{N}$ and each $A^{t}\in\mathbb{R}^{N\times N}$ is a diagonal matrix. Please see Figure˜2 for a visualization.
Then we have $M_{j,i}=\sum_{n=1}^{N}(c_{j})_{n}(A^{j}\cdots A^{i+1})_{n,n}(b_{i})_{n}$ for all $1\leq i\leq j\leq T$.
Note that we separate those terms for different $n_{1},n_{2}\in[N]$. For all $n\in[N]$, let $b^{n},c^{n}\in\mathbb{R}^{T}$ be such that for all $t\in[T]$, $b^{n}_{t}=(b_{t})_{n}$ and $c^{n}_{t}=(c_{t})_{n}$. Please see Figure˜2 for a visualization.
Define $\mathsf{1SS}(\cdot):\mathbb{R}^{T}\rightarrow\mathbb{R}^{T\times T}$ by
$\mathsf{1SS}(a_{1},a_{2},\cdots,a_{T})=\underbrace{M}_{T\times T},$
where for $1\leq t,s\leq T$ it holds that
$M_{t,s}:=\begin{cases}a_{t}\cdots a_{s+1},&\quad\text{for}\quad t\geq s;\\ 0,&\quad\text{for}\quad t<s.\end{cases}$
Then for all $n\in[N]$, we consider the $1$-$S$ masked attention matrix $M^{n}$ as
$M^{n}=\mathsf{1SS}(A^{1}_{n,n},\cdots,A^{T}_{n,n})\odot(b^{n}\cdot c^{n\top}).$
We have $M=\sum_{n=1}^{N}M^{n}$ by simple algebra.
So far we construct the attention-like dual for general SSM. It enables us to further construct the $1$-$S$ masked attention dual for a specific class of diagonal SSM.
4.2 Structured State-Space Duality for Diagonal SSMs with Full-Rank State Matrices
Next we show that, when all the state matrices $A^{t}$ of the state-space model have full rank, the attention-like dual of diagonal SSM turns into $1$-$S$ masked attention dual. Specifically, we use the attention-like representation of $M$ in Section˜4.1 to construct the $1$-$S$ masked attention dual of $M$.
11-$S$ Attention Dual of SSM with Full-Rank Diagonal State Matrices.
Suppose $M\in\mathbb{R}^{T\times T}$ has $N$-SSS representation as in (3.1), where each $A^{t}$ is a diagonal matrix with none-zero determinant. In this case we show that $M$ has a $1$-$S$ masked attention dual.
Specifically, when $\text{det}(A^{t})\neq 0$ for all $t\in[T]$, $M^{n}$ has the representation of (for $n\in[N]$)
$\mathsf{1SS}(1,1,\cdots,1)\odot(b^{\prime n}\cdot c^{\prime n\top}),$
where $b^{\prime n}_{t}:=b^{n}_{t}\cdot(A^{1}_{n,n}\cdots A^{t}_{n,n})$ and $c^{\prime n}_{t}:=c^{n}_{t}/(A^{1}_{n,n}\cdots A^{t}_{n,n})$ for all $t\in[T]$. Then, we let $B^{\prime},C^{\prime}\in\mathbb{R}^{T\times N}$ be such that $B^{\prime}_{:,n}=b^{\prime n}$ and $C^{\prime}_{:,n}=c^{\prime n}$ for all $n\in[N]$, we have
$M=\mathsf{1SS}(1,1,\cdots,1)\odot(B^{\prime}\cdot C^{\prime\top}).$
In this way we construct the $1$-$S$ masked attention dual of diagonal SSM with full-rank state matrices.
4.3 Computational Complexity of Diagonal SSD
We give the concrete computation algorithm of diagonal state-space duality according to the attention-like dual in Section˜4.1 and evaluate its efficiency in three aspects: (i) computation cost; (ii) total memory; (iii) parallelization.
Computation Algorithm of Diagonal SSD.
We define the functions $f:\mathbb{R}^{T}\times\mathbb{R}^{T\times d}\rightarrow\mathbb{R}^{T\times d}$ by $f(x,Y)_{:,s}=x\odot(Y_{:,s})$ for all $s\in[d]$,, and $g:\mathbb{R}^{T}\times\mathbb{R}^{T\times d}\rightarrow\mathbb{R}^{T\times d}$ by $g(x,Y)_{1,:}=Y_{1,:}$ and $g(x,Y)_{t+1,:}=x_{t+1}\cdot g(x,Y)_{t,:}+Y_{t+1,:}$ for $t\in[T-1]$. Consider the SSM layer with state dimension $N$ defined by (3.2), where each $A^{t}$ is a diagonal matrix. The recurrence relation (3.2) also has representation
$\underbrace{Y}_{T\times d}=\underbrace{M}_{T\times T}\cdot\underbrace{X}_{T\times d},$
where
$M_{j,i}=c_{j}^{\top}A^{j}\cdots A^{i+1}b_{i}.$
Express $M$ as $M=\sum_{n=1}^{N}M^{n}$, where $M^{n}=\mathsf{1SS}(A^{1}_{n,n},\cdots,A^{T}_{n,n})\odot(b^{n}\cdot c^{n\top})$. Let $a^{n}\in\mathbb{R}^{T}$ denote $(A^{1}_{n,n},\cdots,A^{T}_{n,n})$ for $n\in[N]$. Denote $Y=M\cdot X$ as $Y=\text{SSM}(X)$. Then $\text{SSM}(X)$ is computed as the following algorithm Algorithm˜1.
Algorithm 1 Diagonal State-Space Dual (SSD)
procedure SSM($X$)
  $Z^{n}\leftarrow f(b^{n},X)\quad\text{for all}\penalty 10000\ n\in[N]$ // Time $\mathcal{O}(NTd)$
  $H^{n}\leftarrow g(a^{n},Z^{n})\quad\text{for all}\penalty 10000\ n\in[N]$ // Time $\mathcal{O}(NTd)$
  $Y^{n}\leftarrow f(c^{n},H^{n})\quad\text{for all}\penalty 10000\ n\in[N]$ // Time $\mathcal{O}(NTd)$
  $Y\leftarrow\sum_{n=1}^{N}Y^{n}$ // Time $\mathcal{O}(NTd)$
  return $Y$
end procedure
Computation Cost Same as Scalar-Identity SSD.
Since each step of Algorithm˜1 takes computation cost of $\mathcal{O}(NTd)$, this algorithm takes total computation cost of $\mathcal{O}(NTd)$ FLOPs. This implies diagonal SSD has the same computation cost as scalar-identity SSD (Dao and Gu, 2024, Theorem 3.8), and matches the maximum parameter complexity for SSD (Dao and Gu, 2024, Appendix A.1).
Total Memory Cost.
The memory cost of the state data $A^{1},\cdots,A^{T},b_{1},\cdots,b_{T},c_{1},\cdots,c_{T}$ is $TN+TN+TN=\mathcal{O}(NT)$. In the first three steps of Algorithm˜1, each step generates $n$ matrices of size $T\times d$, and in the last step of Algorithm˜1, only one matrix of size $T\times d$ is generated. Therefore the memory cost of the intermediate step is $NTd+NTd+NTd+Td=\mathcal{O}(NTd)$. Considering all the memory costs above, we deduce that diagonal state-space duality has total memory cost $\mathcal{O}(NT)+\mathcal{O}(NTd)=\mathcal{O}(NTd)$.
Parallelization.
Note that the first three steps of Algorithm˜1 are operated in parallel for each $n\in[N]$, the diagonal state-space dual has separation into $N$ parallel computation processes, each of them costing time $\mathcal{O}(Td)$. Furthermore, from the definition of $f$ and $g$, note that each column of $X$ operates respectively during the whole processing of Algorithm˜1. Therefore the diagonal state-space dual has further separation into $Nd$ parallel computation processes, each process costing time $\mathcal{O}(T)$.
4.4 General SSMs with 1-SS Masked Attention Duals
We further study the duality between $1$-$S$ masked attention and general SSM, and provide the exact class of general SSM that has $1$-$S$ masked attention dual. Firstly, we state the equivalence between the class of $N$-$S$ matrices and the class of $N$-SSS representable matrices. In particular, we identify a direct 1-1 correspondence between SSMs and $N$-SSS representations. Thus, each SSM has a corresponding $N$-$S$ matrix. Then we study the duality between $1$-$S$ masked attention and general SSM regarding the SSM’s corresponding $N$-$S$ matrix.
Equivalence Between $N$-$S$ Matrices and $N$-SSS Representable Matrices.
We start with a constructive proof for an equivalence introduced in (Dao and Gu, 2024).
Proposition 4.1 (Proposition 3.3 in (Dao and Gu, 2024)).
A lower triangular matrix is $N$-semiseparable iff it is $N$-SSS representable.
Proof.
Please see Section˜A.1 for a detailed proof. ∎
Remark 4.1.
We remark that Proposition˜4.1 complements the proof of (Dao and Gu, 2024, Proposition 3.3). This constructive proof forms a rigorous basis for general SSD and yields an explicit procedure to obtain the $N$-SSS representation from an $N$-$S$ matrix. However, its computational complexity exceeds SSD limit (i.e., Section˜4.3), and fast algorithms for more general SSD are still lacking. We leave the design of such fast algorithms to future work.
Remark 4.2.
The equivalence between $1$-$S$ and $1$-SSS-representable matrices appears in structured-matrix literature such as semiseparable, quasiseparable, and SSS forms (Vandebril et al., 2005). We extend it to $N$-$S$ matrices. Our proof is self-contained and constructive. It follows the causal setting and matches the attention mechanism in transformers. The result is more accessible to the ML community and supports our SSD design.
So far we obtain the equivalence between the class of $N$-$S$ matrices and the class of $N$-SSS-representable matrices. Moreover, we have a direct one-to-one correspondence between $N$-SSS representation and general SSM. Next, we use $N$-$S$ matrices to study the duality between $1$-$S$ masked attention and general SSM.
SSMs with $1$-$S$ Masked Attention Dual.
Here, we provide a necessary and sufficient condition for an SSM to have $1$-$S$ masked attention dual. In particular, we state this condition in terms of the SSM’s corresponding $N$-$S$ matrix. Then, we characterize the possible attention dual for this $N$-$S$ matrix. We start with a few concepts below for convenience of discussion:
Definition 4.1 (Fine $1$-$S$ Matrix).
We say a $1$-$S$ matrix $L=\mathsf{1SS}(a_{1},a_{2},\cdots,a_{t})$ is a fine $1$-$S$ matrix iff $a_{1}a_{2}\cdots a_{t}\neq 0$.
Definition 4.2 (New Column of Lower Triangular Matrix).
Suppose $M\in\mathbb{R}^{T\times T}$ is a lower triangular matrix. We call $M_{t:,t}$ a new column of $M$ iff $M_{t:T+1,t}$ is not in $M_{t:,:t}$’s column space.
We firstly specify the class of $N$-$S$ matrices that has a representation of fine $1$-$S$ masked attention. This prepares us for the later general State-Space Duality in Theorem˜4.1.
Proposition 4.2.
Suppose $M\in\mathbb{R}^{T\times T}$ is an $N$-$S$ lower triangular matrix. Then $M$ has representation of $1$-$S$ masked attention $L\odot(QK^{\top})$ for some $Q,K\in\mathbb{R}^{T\times N}$ and fine $1$-$S$ matrix $L$ iff it has at most $N$ new columns.
Proof.
Please see Section˜A.2 for detailed proof. ∎
Then we deduce the following results from Proposition˜4.2. These results state the necessary and sufficient condition for an SSM to have a general $1$-$S$ masked attention dual.
Lemma 4.1.
Suppose $M\in\mathbb{R}^{T\times T}$ is an $N$-$S$ lower triangular matrix. Then $M$ has representation of $1$-$S$ masked attention $L\odot(QK^{\top})$ for some $Q,K\in\mathbb{R}^{T\times N}$ iff $M$ has several diagonal blocks containing all the none-zero entries of $M$, and each of the diagonal blocks has at most $N$ new columns.
Proof.
The proof consists of two parts.
Part 1: Sufficient Condition.
Here, we show that, if $M$ has several diagonal blocks containing all the none-zero entries of $M$, and each of the diagonal blocks has at most $N$ new columns, then $M$ has the representation of a $1$-$S$ masked attention.
Suppose $M\in\mathbb{R}^{T\times T}$ has $k$ diagonal blocks
$M_{:t_{1},:t_{1}},M_{t_{1}:t_{2},t_{1}:t_{2}},\cdots,M_{t_{k-1}:,t_{k-1}:},\quad\text{for}\quad t_{1},\cdots,t_{k-1},$
such that $1\leq t_{1}<t_{2}<\cdots<t_{k-1}\leq T$. Here we set ${t_{0}}=1$ and $t_{k}=T+1$ for consistency. Then $t_{i}-t_{i-1}$ is the dimension of the $i$-th diagonal block of $M$ for $i\in[k]$.
Denote the $i$-th diagonal block of $M$ as $M^{i}$. Then according to Proposition˜4.2, each diagonal block of $M$ has representation
$M^{i}=\mathsf{1SS}(\underbrace{1,1,\cdots,1}_{t_{i}-t_{1-1}})\odot(Q^{i}K^{i\top})$
for some $Q^{i}.K^{i}\in\mathbb{R}^{(t_{i}-t_{i-1})\times N}$.
Then, letting
$p_{i}:=$ $\penalty 10000\ (0,1,1,\cdots,1)\in\mathbb{R}^{t_{i}-t_{i-1}},$
$p^{\top}:=$ $\penalty 10000\ [p_{1}^{\top},p_{2}^{\top},\cdots,p_{t}^{\top}]\in\mathbb{R}^{1\times T},$
$Q^{\top}:=$ $\penalty 10000\ [Q^{1\top},Q^{2\top},\cdots,Q^{t\top}]\in\mathbb{R}^{n\times T},$
$K^{\top}:=$ $\penalty 10000\ [K^{1\top},K^{2\top},\cdots,K^{t\top}]\in\mathbb{R}^{n\times T},$
we have
$M=\mathsf{1SS}(p)\odot(QK^{\top}).$
Part 2: Necessary Condition.
Here, we show that, if $M$ has the representation of a $1$-$S$ masked attention, then $M$ has several diagonal blocks containing all its entries and each of the diagonal blocks has at most $N$ new columns.
Suppose $M=\mathsf{1SS}(p)\odot(QK^{\top})$ for some $p\in\mathbb{R}^{T}$ and $Q,K\in\mathbb{R}^{T\times N}$. Let k1k-1 be the number of zeros in $p_{2:T+1}\in\mathbb{R}^{T-1}$. Denote the zero positions by $p_{t_{1}}=p_{t_{2}}=\cdots=p_{t_{k-1}}=0$, where $2\leq t_{1}<t_{2}<\cdots<t_{k-1}\leq T$. For all other indices $t\in[T]\setminus\{t_{1},\dots,t_{k-1}\}$, we have $p_{t}\neq 0$.
Then all of $M$’s none-zero elements lie in diagonal blocks
$M_{:t_{1},:t_{1}},M_{t_{1}:t_{2},t_{1}:t_{2}},\cdots,M_{t_{k-1}:,t_{k-1}:}.$
Here, we set $t_{0}=1$ and $t_{k}=T+1$ for consistency. Then, for each of the blocks, we have
$M_{t_{i-1}:t_{i},t_{i-1}:t_{i}}=\mathsf{1SS}(p_{t_{i-1}:t_{i}})\odot(Q_{t_{i-1}:t_{i},:}K_{t_{i-1}:t_{i},:}^{\top})$
for all $i\in[k]$.
Note that the result of $\mathsf{1SS}(p)$ does not depend on the value of $p_{1}$, these diagonal blocks are equivalent to an attention matrix with a fine $1$-$S$ mask. Therefore according to Proposition˜4.2, each of them has at most $N$ new columns.
This completes the proof. ∎
So far we characterize the exact class of $N$-$S$ matrices corresponding to a $1$-$S$ masked attention representation. Given the equivalence between $N$-$S$ matrices and $N$-SSS representable matrices, along with the direct one-to-one correspondence between $N$-SSS-representable matrices and SSMs, we obtain the following theorem:
Theorem 4.1 (General State-Space Duality).
Suppose $M$ is an $N$-$S$ matrix corresponding to an SSM. This SSM has $1$-$S$ masked attention dual iff $M$ has several diagonal blocks containing all the none-zero entries of $M$, and each of the diagonal blocks has at most $N$ new columns.
Remark 4.3.
In Proposition˜4.2, we focus on fine $1$-$S$ masked attention, and in Lemma˜4.1 our conclusion holds for general $1$-$S$ masked attention. In the case of general $1$-$S$ masked attention, $a_{t}$ is possibly 0 for some $t\in[T]$ in the causal mask $L=\mathsf{1SS}(a_{1},a_{2},\cdots,a_{T})$.
5 Limitations of Structured State-Space Duality
In this section, we study the limitations of state-space duality from its two ends: (i) on the attention side, we show the impossibility of extending SSD to practical softmax attention (Vaswani et al., 2017); (ii) on the SSM side, we show that even SSMs with low state dimension or low-rank state matrices may not always admit an SSD extension.
Impossibility of Extending SSD to Softmax Attention.
We present a simple example showing that softmax attention (Vaswani et al., 2017) does not admit state-space duality. Consider a matrix $V\in\mathbb{R}^{T\times T}$ with entries $V_{i,j}=i\times j$ for all $i,j\in[T]$. The matrix $V$ has rank $1$ since every column is a scalar multiple of $(1,2,\dots,T)^{\top}\in\mathbb{R}^{T}$. However, $\mathop{{\rm Softmax}}(V)$ has rank $T$ by the Vandermonde determinant, and in fact, every submatrix of $\mathop{{\rm Softmax}}(V)$ is full rank.
This shows that even when the attention matrix $QK^{\top}$ has very low rank, $\mathop{{\rm Softmax}}(QK^{\top})$ typically expands to full rank $T$. Moreover, any attention matrix with a state-space dual must be $N$-semiseparable, where $N$ is the state dimension of the corresponding SSM. Hence, softmax attention does not have a state-space model dual.
Impossibility of Extending SSD to General SSM with Low State Dimension.
We provide an example for this impossibility in the following proposition. It shows that a general SSM, even with low state dimension or low-rank state matrices, may not admit a $1$-$S$ masked-attention dual.
Proposition 5.1.
Consider the SSM layer with state dimension $N\geq 2$ defined by (3.2), there exist $A^{1:T+1},b_{1:T+1},c_{1:T+1}$ such that the recurrence relation does not have an attention dual.
Proof.
According to Proposition˜4.1, there exist $A^{1:T+1},b_{1:T+1},c_{1:T+1}$ such that the recurrence relation (3.2) has representation $Y=M\cdot X$, where $M:=I_{T\times T}+E^{T,1}$ is a $2$-$S$ matrix. Here
$E^{T,1}_{j,i}:=\begin{cases}1,\quad&\penalty 10000\ j=T\penalty 10000\ \text{and}\penalty 10000\ i=1;\\ 0,\quad&\penalty 10000\ \text{otherwise}.\end{cases}$
We claim that $M$ does not have the representation of $L\odot(QK^{\top})$ where $Q,K\in\mathbb{R}^{T\times N}$ and $L$ is a $1$-$S$ matrix.
Otherwise if $M=L\odot(QK^{\top})$, suppose
$L_{j,i}=\begin{cases}a_{j}\cdots a_{i+1},&\quad\text{for}\quad j\geq i;\\ 0,&\quad\text{for}\quad j<i.\end{cases}$
for $i,j\in[T]$.
Since $M_{T,1}=1$, $L_{T,1}=a_{2}\cdots a_{T}$ is none-zero, i.e. each of $a_{2},a_{3},\cdots,a_{T}$ is none-zero.
Thus, we have
$(QK^{\top})_{i.j}=0,\quad\text{for all}\quad 1\leq j<i\leq T-1.$
[…]
This completes the proof. ∎
Therefore, general SSM does not always have $1$-$S$ masked attention dual, even when the SSM has very low state dimension.
6 Conclusion
Structured state-space models (SSMs) (Dao and Gu, 2024; Gu and Dao, 2023) and attention mechanisms (Vaswani et al., 2017) represent two contrasting paradigms for sequence modeling. SSMs evolve a latent state step by step and run in linear time $O(T)$. Attention mixes all pairs and costs $O(T^{2})$. Structured State-Space Duality (SSD) (Dao and Gu, 2024) provides a theoretical bridge between these two views. We formalize and generalize the structured state-space duality between simple recurrent state-space models (SSMs) and masked attention mechanisms. We extend the duality introduced by (Dao and Gu, 2024) from the scalar-identity case to general diagonal SSMs (Section˜4.1) and show that these models retain the same computational complexity lower bounds while supporting richer dynamics (Section˜4.3). We further characterize the necessary and sufficient condition under which an SSM corresponds to a $1$-semiseparable masked attention mechanism (Section˜4.4). Lastly, we present two negative results (Section˜5): (i) this duality does not extend to standard softmax attention due to a rank explosion in the induced kernel; (ii) even for low-rank SSMs, such duality may still fail at representational level. Together, these results strengthen the theoretical bridge between recurrent SSMs and Transformer-style attention, and broaden the design space for expressive yet efficient sequence modeling.
Acknowledgments
JH would like to thank Alex Reneau, Mimi Gallagher, Sara Sanchez, T.Y. Ball, Dino Feng and Andrew Chen for enlightening discussions on related topics, and the Red Maple Family for support.
JH is partially supported by Ensemble AI and Northwestern University. Han Liu is partially supported by NIH R01LM1372201, NSF AST-2421845, Simons Foundation MPS-AI-00010513, AbbVie, Dolby and Chan Zuckerberg Biohub Chicago Spoke Award.
Appendix
Appendix A Proof of Main Text
A.1 Equivalence between N-SS Matrices and N-SSS-Representable Matrices
Proposition A.1 (Proposition 3.3 in (Dao and Gu, 2024); Proposition˜4.1 Restate).
A lower triangular matrix $M$ is $N$-semiseparable iff it is $N$-SSS-representable, i.e. $M$ being $N$-SSS representable is the necessary and sufficient condition for $M$ to be $N$-semiseparable.
Proof.
Our proof consists of two parts.
Part 1: Sufficient Conditon.
In this part, we take three steps to show that any lower triangular matrix with an $N$-SSS representation is $N$-semiseparable.
  • Step 1: Express $M$ with its $N$-SSS representation. Suppose $M\in\mathbb{R}^{T\times T}$ is a lower triangular matrix with an $N$-SSS representation
    $M_{j,i}=c_{j}^{\top}A_{j}\cdots A_{i+1}b_{i},$ (A.1)
    for all $i,j\in[T]$, where $b_{1},\cdots,b_{T},c_{1},\cdots,c_{T}\in\mathbb{R}^{N}$ and $A^{1},\cdots,A^{T}\in\mathbb{R}^{N\times N}$.
  • Step 2: Express each submatrix $S$ on or below the diagonal of $M$ through the $N$-SSS representation of $M$. Suppose $S$ is a submatrix of $M$ such that each entry of $S$ is on or below the diagonal line of $M$, we have
    $S=M_{j_{1}:j_{2},i_{1}:i_{2}},$
    for some $1\leq j_{1}<j_{2}\leq T+1$, $1\leq i_{1}<i_{2}\leq T+1$ and $i_{2}\leq j_{1}$.
    Denote $S^{1}\in\mathbb{R}^{(j_{2}-j_{1})\times N}$ as
    $S^{1}[:,j]=c_{j+j_{1}-1}^{\top}A^{(j+j_{1}-1)}\cdots A^{j_{1}+1},$
    for $j\in[j_{2}-j_{1}]$.
    Denote $S^{2}\in\mathbb{R}^{N\times(i_{2}-i_{1})}$ as
    $S^{2}[i,:]=A^{j_{1}}\cdots A^{(i+i_{1})}b_{i+i_{1}-1}$
    for $i\in[i_{2}-i_{1}]$.
    Then according to (A.1), we have
    $S=S_{1}\cdot S_{2}.$
  • Step 3: Upperbound the rank of $S$ with $S^{1}$ and $S^{2}$. Since $S^{1}$ and $S^{2}$ both have rank at most $N$, we deduce that $S$ has rank at most $N$.
Part 2: Neccessary Condition.
In this part we take four steps to show that any $N$-semiseparable lower triangular matrix has an $N$-SSS representation. Suppose $M\in\mathbb{R}^{T\times T}$ is an $N$-semiseparable lower triangular matrix.
  • Step 1: Divide $M_{t:,:t+1}$ into the product of two matrices of rank at most $N$. In this step, we represent $M_{t:,:t+1}$ by the product of two low-rank matrices for all $t\in[T]$.
    Since $M$ is $N$-semiseparable, $M_{t:,:t+1}\in\mathbb{R}^{(T-t+1)\times t}$ has rank at most $N$ for each $t\in[T]$. Therefore there exist low-rank matrices $W_{t}\in R^{(T-t+1)\times N}$ and $U_{t}\in\mathbb{R}^{N\times t}$ such that
    $M_{t:,:t+1}=W_{t}\cdot U_{t}.$ (A.2)
    Please see Figure˜4 for a visualization.
    Figure 3: $M_{t:,:t+1}=W^{t}\cdot U^{t}$
    Figure 4: Only the first $r_{t}$ columns (rows) of $W^{t}$ ($U^{t}$) are non-zero.
    Let $r_{t}\leq N$ denote the rank of $M_{t:,:t+1}$ for all $t\in[T]$. Without loss of generality, we construct $W_{t}$ and $U_{t}$ to be such that only the first $r_{t}$ columns of $W_{t}$ are none-zero, and similarly, only the first $r_{t}$ rows of $U_{t}$ are none-zero. Please see Figure˜4 for a visualization.
    This means that the span of $W_{t}$’s column vectors equals the span of $M_{t:T+1,1:t+1}$’s column vectors, and the span of $U^{t}$’s row vectors equals the span of $M_{t:T+1,1:t+1}$’s row vectors.
  • Step 2: Conditions on $A^{t},b_{t},c_{t}$ when $M$ has an $N$-SSS representation. In this step, we characterize the conditions that the state parameters $A^{1:T+1},b_{1:T+1},c_{1:T+1}$ should satisfy when $M$ has an $N$-SSS representation.
    We start by defining $W^{t}$$^{\prime}$ and $U^{t}$$^{\prime}$ for convenience of discussion.
    Set ${W^{t}}^{\prime}=W^{t}_{2:,:}$ and ${U^{t}}^{\prime}=U^{t}_{:,:t}$ for all $t\in[T]$, i.e. $W^{t}$$^{\prime}$ is $W^{t}$ without the first row, and $U^{t}$$^{\prime}$ is $U^{t}$ without the last column. We provide a visualization in Figure˜6.
    Figure 5: ${W^{t}}^{\prime}\cdot U^{t}=W^{t+1}\cdot{U^{t+1}}^{\prime}$
    Figure 6: Set only the first $r_{t}$ columns and the first $r_{t+1}$ rows of $A^{t+1}$ and $A^{t+1}$$^{\prime}$ to be non-zero.
    Then we have
    $W^{t+1}\cdot{U^{t+1}}^{\prime}=M_{t+1:,:t+1}={W^{t}}^{\prime}\cdot U^{t}.$ (A.3)
    Note that if $M$ has an $N$-SSS representation, then for all $t\in[T]$, we have
    $M_{t:,:t+1}=\begin{pmatrix}b_{t}^{\top}\\ b_{t+1}^{\top}A^{t+1}\\ b_{t+2}^{\top}A^{t+2}A^{t+1}\\ \vdots\\ b_{T}^{\top}A^{T}\cdots A^{t+1}\end{pmatrix}\cdot\begin{pmatrix}A^{t}\cdots A^{2}c_{1}&A^{t}\cdots A^{3}c_{2}&\cdots&A^{t}c_{t-1}&c_{t}\end{pmatrix}$. (A.4)
    We expect there exist $A^{1:T+1}$, $b_{1:T+1}$, $c_{1:T+1}$ satisfying
    $\begin{pmatrix}b_{t}^{\top}\\ b_{t+1}^{\top}A^{t+1}\\ b_{t+2}^{\top}A^{t+2}A^{t+1}\\ \vdots\\ b_{T}^{\top}A^{T}\cdots A^{t+1}\end{pmatrix}=W^{t},$ (A.5)
    and
    $\begin{pmatrix}A^{t}\cdots A^{2}c_{1}&A^{t}\cdots A^{3}c_{2}&\cdots&A^{t}c_{t-1}&c_{t}\end{pmatrix}=U^{t},$ (A.6)
    for all $t\in[T]$.
    We observe that (A.5) and (A.6) require ${W^{t}}^{\prime}=W^{t+1}\cdot A^{t+1}$ and ${U^{t+1}}^{\prime}=A^{t+1}\cdot U^{t}$ for all $t\in[T-1]$.
  • Step 3. Verify the existence of $A^{t}$ satisfying the conditions in Step 2. In this step, we show that there exists $A^{t+1}\in\mathbb{R}^{N\times N}$ for all $t\in[T-1]$ satisfying ${W^{t}}^{\prime}=W^{t+1}\cdot A^{t+1}$ and ${U^{t+1}}^{\prime}=A^{t+1}\cdot U^{t}$.
    Since the column space of $W^{t+1}$ equals to the column space of $M_{t+1:,:t+2}$, and the column space of $M_{t+1:,:t+2}$ contains the column space of $W^{t}$$^{\prime}$, we deduce that the column space of $W^{t+1}$ contains the column space of $W^{t}$$^{\prime}$. Thus, there must exist $A^{t+1}\in\mathbb{R}^{N\times N}$ satisfying ${W^{t}}^{\prime}=W^{t+1}\cdot A^{t+1}$.
    For the same reason there exists ${A^{t+1}}^{\prime}\in\mathbb{R}^{N\times N}$ satisfying ${U^{t+1}}^{\prime}={A^{t+1}}^{\prime}\cdot U^{t}$.
[…]
    Now we have
    $W^{t+1}\cdot{A^{t+1}}^{\prime}\cdot U^{t}$ $\penalty 10000\ =W^{t+1}\cdot{U^{t+1}}^{\prime}$
    $\penalty 10000\ =M_{t+1:,:t+1}$
    $\penalty 10000\ ={W^{t}}^{\prime}\cdot U^{t}$
    $\penalty 10000\ =W^{t+1}\cdot A^{t+1}\cdot U^{t},$
    where the 1st step is by ${U^{t+1}}^{\prime}={A^{t+1}}^{\prime}\cdot U^{t}$, the 2nd and 3rd step is by (A.3) (see Figure˜6), and the last step is by ${W^{t}}^{\prime}=W^{t+1}\cdot A^{t+1}$. We provide a visualization in Figure˜7.
    Figure 7: ${W^{t+1}}^{\prime}\cdot{A^{t+1}}^{\prime}\cdot U^{t}={W^{t+1}}^{\prime}\cdot A^{t+1}\cdot U^{t}$
    Since $W^{t+1}$ and $U^{t}$ have rank $r_{t+1}$ and $r_{t}$ perspectively, we deduce that ${A^{t+1}}^{\prime}=A^{t+1}$. So far for any $t\in[T-1]$, we construct $A^{t+1}$ satisfying both ${W^{t}}^{\prime}=W^{t+1}\cdot A^{t+1}$ and ${U^{t}}^{\prime}={A^{t+1}}^{\prime}\cdot U^{t+1}$.
  • Step 4. Construct the $N$-SSS representation of $M$. So far we construct $A^{t}$, $W^{t}$ and $U^{t}$ satisfying (A.2), (A.5), and (A.6). In this step we construct the $N$-SSS representation of $M$ using the $A^{t}$, $W^{t}$ and $U^{t}$.
    For all $t\in[T]$, let $b_{t}$ be the first column of $(W^{t})^{\top}$ and $c_{t}$ be the last column of $U^{t}$. Let $A^{t+1}$ be as constructed above for $t\in[T-1]$ and $A^{1}=I_{N}$. Then
    $M_{j,i}=c_{j}^{\top}A^{j}\cdots A^{i+1}b_{i}.$
    Namely, $M$ has an $N$-SSS representation.
This completes the proof. ∎
A.2 Condition for N-SS Matrix to Have Fine 1-SS Masked Attention Dual
Proposition A.2 (Proposition˜4.2 Restate).
Suppose $M\in\mathbb{R}^{T\times T}$ is an $N$-$S$ lower triangular matrix. Then $M$ has representation of $1$-$S$ masked attention $L\odot(QK^{\top})$ for some $Q,K\in\mathbb{R}^{T\times N}$ and fine $1$-$S$ matrix $L$ iff it has at most $N$ new columns, i.e. $M$ having at most $N$ new columns is the necessary and sufficient condition for $M=L\odot(QK^{\top})$.
Proof.
The proof consists of two parts.
Part 1: Necessary Condition.
In this part we show that $M$ does not have representation of fine $1$-$S$ masked attention if it has more than $N$ new columns.
Suppose
$M=L\odot QK^{\top}$ (A.7)
for some $Q,K\in\mathbb{R}^{T\times N}$ and $L=\mathsf{1SS}(a_{1},a_{2},\cdots,a_{T})$ where $a_{1}a_{2}\cdots a_{T}\neq 0$.
We then multiply the $t$-th row by $\frac{1}{a_{1}a_{2}\cdots a_{t}}$ and multiply the $t$-th column by $a_{1}a_{2}\cdots a_{t}$ for all $t\in[T]$. Note that these operations do not change the number of new columns of $M$.
After these operations, we get a new lower triangular matrix $M^{\prime}$ having at least $N+1$ new columns. We deduce by (A.7) that the lower triangular part of $M^{\prime}$ is exactly the same as the lower triangular part of $QK^{\top}$. Denote $W:=QK^{\top}$, then $W$ has rank at most $N$.
Suppose $M^{\prime}$ has new columns $M^{\prime}_{t_{1}:,t_{1}},M^{\prime}_{t_{2}:,t_{2}},\cdots M^{\prime}_{t_{N+1}:,t_{N+1}}$ for $1\leq t_{1}<t_{2}<\cdots<t_{N+1}\leq T$.
We claim that $W_{:,t_{1}},W_{:,t_{2}},\cdots,W_{:,t_{N+1}}$ are linearly independent. If not so, there exist $c_{1},c_{2},\cdots,c_{N+1}\in\mathbb{R}$ such that at least one of them is none-zero and
$c_{1}W_{:,t_{1}}+c_{2}W_{:,t_{2}}+\cdots+c_{N+1}W_{:,t_{N+1}}=\mathbf{0}_{T}.$
Suppose $n$ is the largest index in $[N+1]$ such that $c_{t_{n}}\neq 0$. Then we have
$c_{1}W_{:,t_{1}}+c_{2}W_{:,t_{2}}+\cdots+c_{n}W_{:,t_{n}}=\mathbf{0}_{T}.$
This implies that
$\penalty 10000\ c_{1}M^{\prime}_{t_{n}:,t_{1}}+c_{2}M^{\prime}_{t_{n}:,t_{2}}+\cdots+c_{n}M^{\prime}_{t_{n}:,t_{n}}$
$=$ $\penalty 10000\ c_{1}W_{t_{n}:,t_{1}}+c_{2}W_{t_{n}:,t_{2}}+\cdots+c_{n}W_{t_{n}:,t_{n}}$
$=$ $\penalty 10000\ \mathbf{0}_{T-t_{n}+1},$
which contradicts to the fact that $M^{\prime}_{t_{n}:,t_{n}}$ is a new column of $M^{\prime}$.
Therefore we deduce that $W_{:,t_{1}},W_{:,t_{2}},\cdots,W_{:,t_{N+1}}$ are linearly independent. This implies that $W$ has rank at least $N+1$, which contradicts to $W=QK^{\top}$. Thus, $M$ does not have the representation of $L\odot(QK^{\top})$, where $Q,K\in\mathbb{R}^{T\times N}$ and $L$ is a fine $1$-$S$ matrix.
Part 2: Sufficient Condition.
In this part we show that any lower triangular matrix with at most $N$ new columns has representation of $L\odot(QK^{\top})$ for some $Q,K\in\mathbb{R}^{T\times N}$ and fine $1$-$S$ matrix $L\in\mathbb{R}^{T\times T}$.
Suppose $M\in\mathbb{R}^{T\times T}$ is a lower triangular matrix having at most $N$ new columns. We now change $M$’s entries above the diagonal to create a matrix with rank at most $N$.
We change the entries column by column from the left to the right.
For $t\in[T]$, if $M_{t:,t}$ is a new column of $M$, remain $M^{\prime}_{:t,t}$ to be $\mathbf{0}_{t-1}$; if $M_{t:,t}$ is not a new column of $M$, there exist $c_{1},c_{2},\cdots,c_{t-1}\in\mathbb{R}$ satisfying
$M_{t:,t}=\sum_{s=1}^{t-1}c_{s}M{t:,s}.$
Set $M^{\prime}_{:t,t}$ to be
$\sum_{s=1}^{t-1}c_{s}M^{\prime}_{:t,s},$
then $M^{\prime}_{:,:t+1}$ and $M^{\prime}_{:,:t}$ have the same column rank.
Given that $M$ has no more than $N$ new columns, we deduce by mathematical induction that $M^{\prime}$ has rank at most $N$. Therefore there exist $Q,K\in\mathbb{R}^{T\times N}$ such that $M^{\prime}=QK^{\top}$.
Then we have
$M=\mathsf{1SS}(1,1,\cdots,1)\odot(QK^{\top}).$
This implies that $M$ has the representation of fine $1$-$S$ masked attention matrix.
This completes the proof. ∎